OpenStudy (anonymous):

Rationalize the denominator of 8/7-√6

OpenStudy (anonymous):

My answer:

OpenStudy (anonymous):

8 sq root 6 over 55

hartnn (hartnn):

49-6=?

OpenStudy (anonymous):

43

OpenStudy (anonymous):

Wait, can I ask how you thought of asking me that?

hartnn (hartnn):

\((a-b)(a+b)=a^2-b^2\)

OpenStudy (anonymous):

56 + 8 square root 6 over 43

hartnn (hartnn):

yeah, thats correct,

hartnn (hartnn):

did u get how ?

hartnn (hartnn):

no?

OpenStudy (anonymous):

Sorry, I was finishing writing the steps down to refer back to it

OpenStudy (anonymous):

I understand how, I just forgot to get the 43

hartnn (hartnn):

ok, good.

OpenStudy (anonymous):

Not only that, but you have to include 8 because that's shown in the problem

OpenStudy (anonymous):

opposed to